a 70% acid solution is to be mixed with a 40% solution to get 90 gallons of a 50% solution. How many gallons of each solution are needed?
Answer by mananth(16946)

Acid
percent ---------------- quantity
acid type I 70.00% ---------------- x gallons
acid type II 40.00% ------ 90 - x gallons
Mixture 50.00% ---------------- 90
Total 90 gallons
70.00% x + 40.00% ( 90 - x ) = 50.00% * 90
70 x + 40 ( 90 - x ) = 4500
70 x + 3600 - 40 x = 4500
70 x - 40 x = 4500 - -3600
30 x = 900
/ 30
x = 30 gallons 70.00% acid type I
60 gallons 40.00% acid type II

m.ananth@hotmail.ca